The following truss is subjected to vertical loads of 20 KN at joints E and D. In addition to the loads, support A settles by 5 mm and member AB and BC are subjected to a temperature drop of 50°C. Given Young’s modulus, E = 200 GPa, cross sectional area for each member, A = 500 mm2 and coefficient of thermal expansion of, α = 1.25 x 10-5/°C.

Find the internal forces in each member using force method.

20 KN 20 KN V E D 3m A 4m 4m
Answer #1

20KN 20KN J20kW j E 3m 5m A To 20 um um RA as redundant assume Let us Hence, According unit load / virtual wook method So-suibiven- So = 5mm = 5x10-3 m A = 500mm 200 kN/mma 1.25x10-5/06 sino = 3/5 5 For Pi 20 KN 20KN D Tore 10 Ayo Ac B calculating meJoint B PBE Ify =0 PBE+PBD sino =0 ☆ -20+PBOX PBD Depec PAB B PBP 100 KN 3 O Efx=0 PBD Coso + PBC -80 KN - SAB=0 ☆ PBC = - 10☆ UAE sino = 1 UAE A UAB ! ų 3 UAE UBE UBET UBD sino =0 Joint A {fy=0 UAE 5 3 sfx=0 => UAE Coso + UAB=0 => UAB = - = 3 x 9 |EFx=0 -> UBOCOS + UBC - UAB = 0 TUBC $=0 wloo Jointo e fy=0 = Yes+ ABD sino =0 UDE RD 2 => UCD - oy - Şx? जय Y CD = -1 UBD LU

Member Pi Ui Li PiUiLi Ui2Li
AB 0 -1.3333 4 0 7.1111111
AE 0 1.66667 5 0 13.888889
BE -20 -1 3 60 3
DE 0 1.33333 4 0 7.1111111
BD 33.3333 1.66667 5 277.7778 13.888889
BC -26.667 -2.6667 4 284.4444 28.444444
CD -40 -1 3 120 3

Hence 0+0+ 60 +0+ 277,776 +284.44 +120 AE РЕ AE R: 5x10-3 AE + (-133X1,25X10°5X50 -2.667x1.25x10*% 50 XY х ч + 7,11 + 13.89 A